Two spherical soap of radius 5 and3 cm are at end of narrow horizontal tube. If surface tension is .03 then pressure difference between2 ends

Grade:11

1 Answers

vanshith
5 years ago
deltap=2T/r
deltap1=2T/r1
deltap2=2T/r2
p2-p1=2T(1/r2-1/r1)
         =2×0.03×(0.05-0.03/0.05×0.03)
deltap=4/5=0.8